# Separability from Multipartite Measures

### 03 May 2026 v1

@@ 1 Introduction
The Positive Partial Transposition (PPT) criterion provides a powerful tool for detecting quantum entanglement in bipartite systems [35]. For a density matrix $\rho$, one considers its partial transpose. If $\rho$ is separable, then all eigenvalues of its partial transpose are nonnegative [35]. Therefore, the presence of any negative eigenvalue signals quantum entanglement [35]. This idea underlies negativity [44] and logarithmic negativity [36], which are widely used entanglement measures [42, 45]:
@@
Moreover, negativity and logarithmic negativity have several practical advantages: they are computable for mixed states, both analytically and numerically [11], without requiring a convex-roof construction [36], and they are relatively accessible in experiments [41]. These features make it especially useful to study the Rényi negativities [13, 50], whose spectra encode richer information about entanglement. However, the PPT criterion has an important limitation: It is necessary and sufficient only in low dimensions, specifically for $2\times 2$ and $2\times 3$ Hilbert spaces [26]. In higher dimensions, there exist entangled states that are PPT but still entangled [26]. Thus, PPT is no longer a sufficient criterion for separability in general.
@@
Eq. (10) reveals several key features:
  • Permutation symmetry. The expression (10) is manifestly symmetric under $A$, $B$, $C$, which is essential for a genuine multipartite measure.
  • Decomposition of entanglement structure. The measure separates into:
    • $I_{5}$: a global, higher-order correlation (experimentally accessible),
    • pairwise tangles: bipartite entanglement content,
    • $\tau_{ABC}$: genuine tripartite entanglement.
  • Physical interpretation. $\phi_{ABC}$ effectively combines:
    • detectability (via $I_{5}$),
    • monotonicity (through its construction),
    • completeness (capturing both bipartite and tripartite structures).
Eq. (10) shows that although $I_{5}$ alone is not sufficient, it becomes part of a well-defined entanglement measure when combined with lower-order measures. In this sense, $\phi_{ABC}$ provides a bridge between experimentally measurable quantities and theoretically well-defined entanglement measures. More importantly, it supports our broader goal of constructing permutation-symmetric quantities that can diagnose multipartite separability and entanglement, which we now discuss in the general multipartite setting.
@@
The difficulty of classifying multipartite entanglement comes from the fact that the number of inequivalent entanglement classes grows combinatorially with qubit number. This complexity already appears in the simplest nontrivial cases: three qubits admit 2 inequivalent genuine entanglement classes [14], while four qubits admit 9 inequivalent families [43]. Thus, unlike bipartite systems, where quantities such as the entanglement entropy provide a complete ordering for pure states, multipartite entanglement cannot be totally ordered or fully quantified by a single scalar, but by a hierarchy of structures.
@@
which exhibits purely global entanglement: after tracing out one or more qubits, all reduced density matrices become separable, and no local entanglement survives [39]. More remarkably, the converse direction also holds in an appropriate sense: detecting the separability of all reduced subsystems effectively identifies a GHZ-type structure. In local quantum field theory (QFT), the situation is even more extreme. The theory has infinitely many degrees of freedom, and its entanglement structure is further complicated by ultraviolet (UV) divergences that require regularization. Moreover, entanglement is generically present in spatial subregions [47]. Therefore, in local QFT, separability in position space is essentially lost. While position-space separability is lost in local QFT, it can re-emerge in momentum space. For example, the vacuum state in Lifshitz theory can exhibit mode-wise separability [2].
To make progress, one usually combines the replica trick [25], twist operators [6], and conformal field theory (CFT) [18, 28]. This framework provides a systematic way to compute Rényi entropy and Rényi negativity [5]. Conformal symmetry fixes the universal part of these quantities. At the same time, theory-dependent data, such as operator product expansion (OPE) coefficients [18, 28], can be obtained from anti-de Sitter/conformal field theory (AdS/CFT) correspondence [29, 22, 46, 10].
The AdS/CFT correspondence is a conjectured relationship between two distinct classes of physical theories. On one side lies Anti-de Sitter (AdS) space, which arises in theories of quantum gravity. In contrast, on the other side lie conformal field theories (CFTs), which are QFTs defined on the asymptotic boundary of spacetime. This idea is deeply motivated by insights from black hole thermodynamics [23], particularly the observation that the number of degrees of freedom associated with a region containing a black hole scales with its horizon area rather than its volume. This insight led to the broader holographic principle and, ultimately, to the AdS/CFT correspondence.
Although a complete theory of quantum gravity remains elusive, the AdS/CFT framework provides a powerful tool for studying strongly coupled quantum systems. In particular, it enables the exploration of semiclassical gravity and the large central-charge CFTs via the holographic dictionary [46]. In many cases, computations on the gravitational (AdS) side are technically simpler than direct calculations in the CFT, allowing one to probe regimes of QFT that are otherwise analytically inaccessible. The intrinsic difficulty of CFT and more general QFTs largely stems from the fact that they are defined on infinite-dimensional Hilbert spaces. The AdS/CFT correspondence offers a way to reorganize this complexity into a geometric framework, thereby providing new insights into strongly correlated quantum systems.
We can also use qudit systems as a controlled bridge to QFT. This works because the qudit systems mimic QFT subregions from the large local dimension. Therefore, they provide a clean laboratory for multipartite structures [19, 20] and separability criteria.
@@ 1.1 Summary of Results
  • The four-qubit pure state is the simplest nontrivial system whose one-qubit reductions yield tripartite mixed states, making it the natural arena for systematically applying the separability criterion. We identify a complete set of eighteen independent entanglement measures for four-qubit pure states, summarized in Table 1. Among them, $\phi_{ABC}$ and its permutations play a distinguished role: they are closely related to $I_{5}$ and provide necessary and sufficient criteria for full separability of tripartite states. Furthermore, we show that the remaining quadripartite measures— $4|H|^{2}$, $|\Sigma|^{1/2}$, and $|\Pi|^{1/3}$—are all generated by the hyperdeterminant $\Delta$, motivating the search for new entanglement measures derived from $\Delta$.
@@ 2.1 Tripartite Pure State
The lower bound is $I_{5}\geq 2/9$, which saturates when the state is W-state. See Appendix A for a proof [^3] Ref. [33] stated that the minimum of $I_{5}$ is attained by the W-state, with the value $2/9$. We have not found a proof of this statement in the existing literature.. Therefore, by measuring $I_{5}$, one can determine how close a 3-qubit pure state is to the W-state.
@@ 2.2 Three-Qubit Mixed State
We now show that $\phi_{ABC}$ can serve as a necessary and sufficient criterion for full separability of three-qubit mixed states.
@@
One the one hand, since $\tau_{A|B}$, $\tau_{B|C}$, $\tau_{C|A}$, $\tau_{ABC}$ and $1-I_{5}$ are all nonnegative, $\phi_{ABC}=0$ implies $1-I_{5}=0$. According to Proposition 1, since $1-I_{5}$ is a continuous nonnegative function on the pure-state manifold, there exists a pure-state decomposition $\rho=\sum_{j}p_{j}|\psi_{j}\rangle\langle\psi_{j}|$ such that
@@ 3 Entanglement Measures for Four-Qubit Pure States
Four-qubit pure states provide the simplest but nontrivial setting in which the criteria in Sec. 2 becomes directly useful: tracing out one qubit produces a tripartite mixed state. At the same time, the pure-state system itself already exhibits a rich multipartite structure. A four-qubit pure state modulo local unitary transformations is characterized by eighteen independent real parameters [8], thus a complete description should involve eighteen independent quantities. We identify such a set explicitly in this section. As summarized in Table 1, these eighteen measures consist of six bipartite measures, eight tripartite measures, and four quadripartite measures. We now introduce them and explain how together they form a complete set of entanglement measures for four-qubit pure states. In particular, we introduce how the hyperdeterminant generate these entanglement measures. Finally, we evaluate these entanglement measures for representative states and list the nine entangled ways for 4-qubit pure states.
@@ 3.3 Quadripartite Entanglement Measures
then like Eq. (43), the four-tangle $4|H|^{2}$ can be computed from the squart roots of $\rho_{ABCD}\tilde{\rho}_{ABCD}$ eigenvalues [40, 48]
@@ 5.2 Subsystem Separability and Generalized GHZ-States
Because the local QFT cannot have a separable state for the spatial partition, the generalization of the qudit-state proof already implies that the local QFT for the spatial partition does not allow a GHZ-type state if the qudit state can approach the QFT. The GHZ-type state can only explore the mode-wise entanglement, such as in Lifshitz’s vacuum state [2]. We will demonstrate the computability via third-order negativity in CFT2 in the next section.
@@ 6 Application in CFT
where the permutation elements $g_{j}\in S_{3}$ are determined by the gluing conditions in the replica manifold. For the third-order negativity, each insertion is associated with a $3$-cycle, so the conformal dimensions of the three twist operators are identical,

### 11 May 2026 v2

@@ 1 Introduction
The Positive Partial Transposition (PPT) criterion provides a powerful tool for detecting quantum entanglement in bipartite systems [35]. For a density matrix $\rho$, one considers its partial transpose. If $\rho$ is separable, then all eigenvalues of its partial transpose are nonnegative [35]. Therefore, the presence of any negative eigenvalue signals quantum entanglement [35]. This idea underlies negativity [44] and logarithmic negativity [36], which are widely-used entanglement measures [42, 45]:
@@
Moreover, negativity and logarithmic negativity have several practical advantages: they are computable for mixed states, both analytically and numerically [11], without requiring a convex-roof construction [36], and they are relatively accessible in experiments [41]. These features make it especially useful to study the Rényi negativities [13, 50], whose spectra encode richer information about entanglement. However, the PPT criterion has a limitation: It is necessary and sufficient only in low dimensions, specifically for $2\times 2$ and $2\times 3$ Hilbert spaces [25]. In higher dimensions, there exist entangled states that are PPT but still entangled [25]. Thus, PPT is no longer a sufficient criterion for separability in general.
@@
Eq. (10) exhibits the manifest permutation symmetry among $A$, $B$, and $C$. Furthermore, it shows that although $I_{5}$ alone is not sufficient, it becomes part of a well-defined entanglement measure oncecombined with specific bipartite and tripartite measures. In this sense, $\phi_{ABC}$ provides a bridge between experimentally measurable quantities ($I_{5}$) and theoretically well-defined entanglement measures. More importantly, it supports our broader goal of constructing permutation-symmetric quantities that can diagnose multipartite separability and entanglement, which we now discuss in the general multipartite setting.
@@
The difficulty of classifying multipartite entanglement comes from the fact that the number of inequivalent entanglement classes grows combinatorially with qubit number. This complexity already appears in the simplest nontrivial cases: three qubits admit 2 inequivalent genuine entanglement classes [14], while four qubits admit 9 inequivalent families [43]. Thus, unlike bipartite systems, where quantities such as the entanglement entropy provide a complete ordering for pure states, multipartite entanglement cannot be totally ordered or fully quantified by a single scalar, but necessarily by a hierarchy of structures.
@@
which exhibits purely global entanglement: after tracing out one or more qubits, all reduced density matrices become separable, and no local entanglement survives [39]. More remarkably, the converse direction also holds in an appropriate sense: detecting the separability of all reduced subsystems effectively identifies a GHZ-type structure. In this paper, we will show that the above relation between subsystem separability and GHZ-type states also holds for multipartite qudit systems.
This observation motivates a broader question: how far can separability-based characterizations of multipartite entanglement be extended beyond finite-dimensional systems? Quantum field theory (QFT) provides a natural but much more difficult arena for this question. The theory has infinitely many degrees of freedom, and its entanglement structure is further complicated by ultraviolet (UV) divergences that require regularization. Moreover, entanglement is generically present in spatial subregions [47]. Therefore, in local QFT, separability in position space is essentially lost, but it can re-emerge in momentum space. For example, the vacuum state in Lifshitz theory can exhibit mode-wise separability [2].
To study such infinite-dimensional systems, one usually combines the replica trick [24], twist operators [6], and conformal field theory (CFT) [18, 28]. This framework provides a systematic way to compute Rényi entropy and Rényi negativity [5]. Conformal symmetry fixes the universal part of these quantities. At the same time, theory-dependent data, such as operator product expansion (OPE) coefficients [18, 28], can be obtained from anti-de Sitter/conformal field theory (AdS/CFT) correspondence [29, 22, 46, 26, 10]. We can also use qudit systems as a controlled bridge to QFT. This works because the qudit systems mimic QFT subregions from the large local dimension. Therefore, they provide a clean laboratory for multipartite structures [19, 20] and separability criteria.
@@ 1.1 Summary of Results
  • The four-qubit pure state is the simplest nontrivial system whose one-qubit reductions yield tripartite mixed states, making it the natural arena for systematically applying the separability criterion. We identify a complete set of eighteen independent entanglement measures for four-qubit pure states, as summarized in Table 1. Among them, $\phi_{ABC}$ and its permutations play a distinguished role: they are closely related to $I_{5}$ and provide necessary and sufficient criteria for full separability of tripartite states. Furthermore, we show that the remaining quadripartite measures— $4|H|^{2}$, $|\Sigma|^{1/2}$, and $|\Pi|^{1/3}$—are all generated by the hyperdeterminant $\Delta$, motivating the search for new entanglement measures derived from $\Delta$.
@@ 2.1 Tripartite Pure State
The lower bound is $I_{5}\geq 2/9$, which saturates when the state is W-state [33]. See Appendix A for a detailed proof. Therefore, by measuring $I_{5}$, one can determine how close a 3-qubit pure state is to the W-state.
@@ 2.2 Three-Qubit Mixed State
We then show that $\phi_{ABC}$ can serve as a necessary and sufficient criterion for full separability of three-qubit mixed states.
@@
On the one hand, since $\tau_{A|B}$, $\tau_{B|C}$, $\tau_{C|A}$, $\tau_{ABC}$ and $1-I_{5}$ are all nonnegative, $\phi_{ABC}=0$ implies $1-I_{5}=0$. According to Proposition 1, since $1-I_{5}$ is a continuous nonnegative function on the pure-state manifold, there exists a pure-state decomposition $\rho=\sum_{j}p_{j}|\psi_{j}\rangle\langle\psi_{j}|$ such that
@@ 3 Entanglement Measures for Four-Qubit Pure States
Four-qubit pure states provide the simplest but nontrivial setting in which the criteria in Sec. 2 become directly useful: tracing out one qubit produces a tripartite mixed state. At the same time, the pure-state system itself already exhibits a rich multipartite structure. A four-qubit pure state modulo local unitary transformations is characterized by eighteen independent real parameters [8], thus a complete description should involve eighteen independent quantities. We identify such a set explicitly in this section. As summarized in Table 1, these eighteen measures consist of six bipartite measures, eight tripartite measures, and four quadripartite measures. We now introduce them and explain how together they form a complete set of entanglement measures for four-qubit pure states. In particular, we introduce how the hyperdeterminant generates these entanglement measures. Finally, we evaluate these entanglement measures for representative states and list the nine entangled ways for 4-qubit pure states.
@@ 3.3 Quadripartite Entanglement Measures
then like Eq. (43), the four-tangle $4|H|^{2}$ can be computed from the square roots of $\rho_{ABCD}\tilde{\rho}_{ABCD}$ eigenvalues [40, 48]
@@ 5.2 Subsystem Separability and Generalized GHZ-States
Because the local QFT cannot have a separable state for the spatial partition, the generalization of the qudit state proof already implies that the local QFT for the spatial partition does not allow a GHZ-type state if the qudit state can approach the QFT. The GHZ-type state can only explore the mode-wise entanglement, such as in Lifshitz’s vacuum state [2]. We will demonstrate the computability via third-order negativity in CFT2 in the next section.
@@ 6 Application in CFT
where the permutation elements $g_{j}\in S_{3}$ are determined by the gluing conditions in the replica manifold. For the third-order negativity, each insertion is associated with a $3$-cycle, so the conformal dimensions of the three twist operators are identical [26],